QUESTION ONE (30 Marks)


Consider a closed economy that is characterized by the following equations:


Y = C + I + G


C = 900 + 0.5(Y − T)


I = 750 − 30r


T = 800


G = 1200


Md = Ms



Ms = 1500


Mt = 0.7Y


Msp = −80r


Where Y is the GDP, C is private consumption expenditure, I is the Investment expenditure, G


is government expenditure, T is tax revenues, Ms


is money supply, Mt


is transaction demand


for money, Msp is the speculative demand for money and r is the interest rate (in % points).


a) Derive (Md⁄P) the demand for real money balances equation (where P is the aggregate price level

1
Expert's answer
2022-03-15T16:50:53-0400

Money demand= Mt + Msp= 0.7Y+ -80r

= 0.7Y- 80r

LM = Md = Ms

0.7Y- 80r= 1500

0.7Y= 1500+ 80r

Y= 2142.86+ 114.286r